# Conrol-Translated Finsler-type structure and Anisotropic Ginzbor–Landau models

### v1

@@ Abstract
This paper develops a geometric and analytical extension of the Finsler–Ginzburg–Landau framework by introducing a distributed control field acting as a translation in the tangent bundle. Within this formulation, the classical Tonelli Lagrangian is deformed into a control–translated Finsler structure, whose Legendre dual induces a uniformly elliptic operator and a convex energy functional preserving the essential variational features of the anisotropic model. This approach provides a rigorous analytical setting for coupling external control fields with the intrinsic Finsler geometry of anisotropic superconductors. The study establishes the convexity, coercivity, and regularity properties of the induced energy functional and proves the existence of controlled minimizers through variational arguments on admissible configurations. In the asymptotic regime as the Ginzburg–Landau parameter tends to zero, a detailed $\Gamma$–convergence analysis yields a renormalized energy $W_{u}$ governing vortex interactions under control translation, quantifying the modification of the Green kernel and the self-energy due to the field $u(x)$. The results demonstrate that the control translation preserves the underlying Finsler structure while introducing a new geometric degree of freedom for manipulating and stabilizing vortex configurations.
@@
Keywords: Finsler geometry; anisotropic Ginzburg–Landau; vortices; renormalized energy; $\Gamma$-convergence; control field; Tonelli Lagrangian.
[2020] Mathematics Subject Classification: 53B40, 35Q56, 49J45, 35B40, 82D55, 49K20
@@ 1 Introduction
The variational structure of the Ginzburg–Landau theory has long served as a fundamental framework for understanding the formation and interaction of vortices in superconductivity and related condensed-matter systems. In its classical Euclidean formulation, the model relies on an isotropic quadratic energy density involving the norm $|y|$ on tangent spaces, which leads to elliptic equations with rotationally invariant coefficients and vortex configurations constrained by isotropic symmetry. Yet in many modern materials—notably those exhibiting crystalline anisotropy, layered structures, or nematic order—the isotropic approximation fails to capture the essential geometric and energetic features of the system. This motivates a deeper geometric reformulation of the Ginzburg–Landau functional within the broader setting of Finsler geometry.
@@
Finsler geometry provides a natural analytical and geometric extension of the Riemannian framework by allowing the norm to depend smoothly on both position and direction. The resulting class of strongly convex functions $F(x,y)$ yields anisotropic Laplacians, non-Euclidean metric structures, and curvature-dependent elliptic operators that are particularly well suited for modeling anisotropic physical media [11, 3, 8, 12]. See also recent analyses of vortex stability in anisotropic Ginzburg–Landau systems in [10]. In [1], the author introduced a Finslerian formulation of the Ginzburg–Landau theory, proving the existence of anisotropic vortices, deriving the associated Euler–Lagrange equations, and characterizing the renormalized interaction energy governed by the Finsler Laplacian. That work demonstrated that the anisotropic energy landscape can be rigorously captured within a Finsler geometric setting, providing a coherent mathematical model for superconductors with direction-dependent coherence lengths. Subsequently, in [2], the analysis was extended to the asymptotic regime of vortex concentration, where the renormalized energy $W_{F}$ emerges as the $\Gamma$–limit of the Finsler–Ginzburg–Landau functional, linking the geometry of the base manifold to the interaction laws of vortices [9].
@@
The present paper advances this geometric framework by introducing a distributed control field $u(x)$ acting as a translation in the tangent bundle, leading to a control–translated Tonelli–Finsler structure. In Section 2, we rigorously define the Lagrangian $\varphi_{u}(x,y)$ and its convex dual $\varphi_{u}^{*}(x,\xi)$, together with the associated energy functional $E_{u}[\psi,A]$ and the elliptic operator $\Delta_{u}$. Their explicit analytical forms are provided in Eqs. (2)–(7), where the structural properties of convexity, coercivity, and ellipticity are established in detail. These constructions serve as the geometric and variational foundation of the Finsler–Ginzburg–Landau model under control translation, establishing a consistent link between anisotropic geometry and control theory.
@@
Building on this foundation, the paper develops the analytical framework of the control–translated Tonelli–Finsler energy and its asymptotic behavior. Through a detailed $\Gamma$–convergence analysis as $\varepsilon\to 0$, the renormalized energy $W_{u}$ is derived, governing vortex interactions in the presence of control and describing how $u(x)$ modifies both the Green kernel and the self-interaction potential. The final sections establish a variational and dynamical formulation, including the Euler–Lagrange system, the Hamiltonian structure of vortex motion, and the associated optimal control problem minimizing $E_{u}$.
@@
The results presented here extend the geometric program initiated in [1, 2] by embedding control theory directly into the convex–geometric formulation of the Finsler–Ginzburg–Landau model. The novelty of this work lies in the introduction of a control–translated Tonelli structure that couples Finsler geometry, convex analysis, and optimal control in a unified variational framework. This synthesis between Tonelli dynamics, Finsler geometry, and control analysis yields a mathematically consistent and geometrically interpretable mechanism for the stabilization, manipulation, and optimization of vortex structures in anisotropic superconducting media.
2 Control-–Translated Tonelli Structure and Deformed Energy Functional
@@
remains smooth, strictly convex, and superlinear in $y$. Hence, it belongs to the class of Tonelli Lagrangians, that is, $C^{2}$ functions $L:TM\to\mathbb{R}$ which are strictly convex and superlinear in the fiber variable (cf. [7]). Tonelli Lagrangians generalize the analytic structure of Finsler geometry without assuming homogeneity, while preserving Legendre duality and Hamiltonian regularity. They provide a natural framework for control–-dependent deformations where translation in the fiber variable encodes a geometric control field.
Let $(M,F)$ be a smooth compact manifold endowed with a strongly convex Finsler function $F:TM\to[0,\infty)$. Denote by $F^{*}(x,\xi)$ its convex dual. For a smooth control field $u\in C^{1}(M,TM)$ we define the control–translated Tonelli Lagrangian $\phi_{u}$ as in (2), and its convex conjugate on $T_{x}^{*}M$ by
@@
Let $\psi:M\to\mathbb{C}$ denote the order parameter and $A\in\Omega^{1}(M)$ a magnetic potential. Define the covariant derivative $D_{A}\psi=(d-iA)\psi$. The energy functional induced by the control–translated Tonelli structure is
@@
where $d\mu$ denotes the Busemann–-Hausdorff volume form associated with the base Finsler structure $F$. Since the translation $y\mapsto y-u(x)$ preserves the volume element on each tangent fiber, $d\mu$ remains invariant under the control deformation. This choice ensures that the energy is geometrically well-defined and consistent with the underlying Tonelli structure (cf. [3]).
@@
Substituting (4) with $\xi=D_{A}\psi$ into (2) yields (9). The first term controls $\|D_{A}\psi\|_{L^{2}}$ by equivalence of Finsler and background norms. Using the Cauchy–Schwarz and Young inequalities,
@@
The Euler–-Lagrange equations corresponding to $\mathcal{E}_{u}$ capture how the control interacts with the gauge and order fields through the geometric translation.
@@
Let $u_{k}\to u$ in $L^{\infty}(M;TM)$. Then $\mathcal{E}_{u_{k}}$ $\Gamma$–converges to $\mathcal{E}_{u}$ with respect to weak convergence in $H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$. Consequently, minimizers and gradient flows depend continuously on $u$ (cf. [6]).
@@
From (9), the uu–-dependence is affine and continuous in $L^{\infty}$. The $\Gamma$–-liminf inequality follows from convexity and weak lower semicontinuity of $F^{*}(x,\cdot)^{2}$. A recovery sequence is given by $(\psi_{k},A_{k})=(\psi,A)$ because $\int_{M}\langle D_{A}\psi,u_{k}-u\rangle\,d\mu\to 0$. Therefore $\mathcal{E}_{u_{k}}\to\mathcal{E}_{u}$ in the $\Gamma$–sense. ∎
The functional $\mathcal{E}_{u}$ thus arises rigorously from the control–translated Tonelli Lagrangian $\phi_{u}$. The additional term $\langle D_{A}\psi,u\rangle$ is not a phenomenological assumption but an exact analytical consequence of convex duality. In this setting, control acts geometrically as a translation in the fiber variable, providing a mathematically consistent mechanism coupling geometry, control, and anisotropic superconductivity.
3 Stability and Renormalized Energy in
Control–Translated Tonelli Structures
In this section we examine the asymptotic regime of the control–translated Tonelli Ginzburg–Landau functional $\mathcal{E}_{u}$ introduced previously, focusing on the emergence of renormalized energy, vortex concentration, and stability of critical configurations under control perturbations. The analysis combines the convex–analytic structure of $\phi_{u}^{*}$ with classical $\Gamma$–convergence arguments and vortex theory in anisotropic settings [4, 13].
Let $M$ be a compact oriented two–dimensional manifold equipped with the Tonelli structure $(F_{u},F_{u}^{*})$ as defined in Section 2. For $\varepsilon>0$ sufficiently small, consider configurations $(\psi,A)\in H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$ minimizing
@@ 3 Stability and Renormalized Energy in
The measure $d\mu$ is the Busemann–Hausdorff volume of the base Finsler structure $F$, invariant under the translation $y\mapsto y-u(x)$ as explained before. Denote by $j_{A}=\Im(\bar{\psi}D_{A}\psi)$ the supercurrent and by $\rho=|\psi|^{2}$ the density.
@@
By the compactness and energy quantization principles of Bethuel–Brezis–Hélein [4] and the anisotropic ball construction and Jacobian estimates of Sandier–Serfaty [13], one deduces $\rho_{\varepsilon}\to 1$ in $L^{1}(M)$ and hence in measure. Writing $\psi_{\varepsilon}=\rho_{\varepsilon}^{1/2}e^{i\theta_{\varepsilon}}$, the current decomposes as $j_{A_{\varepsilon}}=\rho_{\varepsilon}(D\theta_{\varepsilon}-A_{\varepsilon})$. The coarea formula and the construction of vortex balls as in [4, 13] yield a finite family of disjoint balls $\{B(a_{i}^{\varepsilon},r_{i}^{\varepsilon})\}$ covering the defect set with $\sum_{i}r_{i}^{\varepsilon}\to 0$ and
@@
normalized by $\int_{M}G_{u}(x,y)\,d\mu(x)=0$. By the ellipticity of $\Delta_{u}$ (Theorem 2.2) and compactness of $M$, $G_{u}$ exists, is symmetric, and belongs to $C^{2,\alpha}_{\mathrm{loc}}(M\times M\setminus\mathrm{diag})$; this follows from Lax–Milgram, Fredholm alternative, and Schaüder theory [5], with the nonlinearity absorbed in the definition of $\Delta_{u}$ via the convex dual $\phi_{u}^{*}$ and the uniform positive definiteness [11].
@@
The function $\Phi_{u}$ encodes the effect of control translation on the self–interaction potential.
@@
Here we used the defining relation $\Delta_{u}G_{u}(\cdot,y)=\delta_{y}-1/\mu(M)$ and the fact that $\mathrm{curl}\,\mathcal{J}_{\varepsilon}=2\pi\sum_{i}d_{i}\delta_{a_{i}^{\varepsilon}}$ outside core balls. Passing to the limit along the convergent subsequence for $\nu_{\varepsilon}$ gives (16)–(17). The specific affine contribution (18) comes from the term $\int\langle\mathcal{J}_{\varepsilon},u\rangle d\mu$ via the identity
@@
… which follows from the Green representation and Schauder theory for uniformly elliptic operators (see [5], and cf. [4] for the GL context]).
@@
In the limit $\varepsilon\to 0$, the supercurrent $j_{A}$ converges weakly to a measure–valued current $J_{u}$ satisfying $\mathrm{curl}\,J_{u}=2\pi\sum_{i}d_{i}\delta_{a_{i}}$ and
@@
The positions $a_{i}$ are local minima of $\mathcal{W}_{u}$, and under small perturbations of $u$ they vary smoothly according to
@@
Suppose $u$ is a smooth control field with small $\|u\|_{C^{1}}$. Then the equilibrium configuration of vortices minimizing $\mathcal{W}_{u}$ differs from the unperturbed lattice of the isotropic Finsler–Ginzburg–Landau model by an affine deformation of order $O(\|u\|_{C^{1}})$. Moreover, the corresponding minimal energy satisfies
@@
4 Variational Control and Vortex
Dynamics in Tonelli–Finsler Geometry
The variational formulation developed in the previous sections naturally induces a dynamical framework where the control field $u(x)$ interacts with the geometric flow of vortices. We now derive and analyse the Euler–Lagrange equations of the control–translated Tonelli–Finsler energy $\mathcal{E}_{u}$ and describe the induced dynamics of vortex centers in the anisotropic setting.

### v2

# Control-Translated Finsler-type structure and Anisotropic Ginzbor-Landau models

@@ Abstract
This paper develops a geometric and analytical extension of the Finsler-Ginzburg-Landau framework by introducing a distributed control field acting as a translation in the tangent bundle. Within this formulation, the classical Tonelli Lagrangian is deformed into a control-translated Finsler structure, whose Legendre dual induces a uniformly elliptic operator and a convex energy functional preserving the essential variational features of the anisotropic model. This approach provides a rigorous analytical setting for coupling external control fields with the intrinsic Finsler geometry of anisotropic superconductors. The study establishes the convexity, coercivity, and regularity properties of the induced energy functional and proves the existence of controlled minimizers through variational arguments on admissible configurations. In the asymptotic regime as the Ginzburg-Landau parameter tends to zero, a detailed $\Gamma$-convergence analysis yields a renormalized energy $W_{u}$ governing vortex interactions under control translation, quantifying the modification of the Green kernel and the self-energy due to the field $u(x)$. The results demonstrate that the control translation preserves the underlying Finsler structure while introducing a new geometric degree of freedom for manipulating and stabilizing vortex configurations.
@@
[…]
[2020] Mathematics Subject Classification: Primary: 53B40, 49J45, 35Q56; Secondary: 49S05, 58B20, 35A15, 35B40, 82D55.
@@ 1 Introduction
The variational structure of the Ginzburg-Landau theory has long served as a fundamental framework for understanding the formation and interaction of vortices in superconductivity and related condensed-matter systems. In its classical Euclidean formulation, the model relies on an isotropic quadratic energy density involving the norm $|y|$ on tangent spaces, which leads to elliptic equations with rotationally invariant coefficients and vortex configurations constrained by isotropic symmetry. Yet in many modern materials-notably those exhibiting crystalline anisotropy, layered structures, or nematic order-the isotropic approximation fails to capture the essential geometric and energetic features of the system. This motivates a deeper geometric reformulation of the Ginzburg-Landau functional within the broader setting of Finsler geometry.
@@
Finsler geometry provides a natural analytical and geometric extension of the Riemannian framework by allowing the norm to depend smoothly on both position and direction. The resulting class of strongly convex functions $F(x,y)$ yields anisotropic Laplacians, non-Euclidean metric structures, and curvature-dependent elliptic operators that are particularly well suited for modeling anisotropic physical media [11, 3, 8, 12]. See also recent analyses of vortex stability in anisotropic Ginzburg-Landau systems in [10]. In [1], the author introduced a Finslerian formulation of the Ginzburg-Landau theory, proving the existence of anisotropic vortices, deriving the associated Euler-Lagrange equations, and characterizing the renormalized interaction energy governed by the Finsler-Laplacian. That work demonstrated that the anisotropic energy landscape can be rigorously captured within a Finsler geometric setting, providing a coherent mathematical model for superconductors with direction-dependent coherence lengths. Subsequently, in [2], the analysis was extended to the asymptotic regime of vortex concentration, where the renormalized energy $W_{F}$ emerges as the $\Gamma$–limit of the Finsler-Ginzburg-Landau functional, linking the geometry of the base manifold to the interaction laws of vortices [9].
@@
The present paper advances this geometric framework by introducing a distributed control field $u(x)$ acting as a translation in the tangent bundle, leading to a control-translated Tonelli-Finsler structure. In Section 2, we rigorously define the Lagrangian $\varphi_{u}(x,y)$ and its convex dual $\varphi_{u}^{*}(x,\xi)$, together with the associated energy functional $E_{u}[\psi,A]$ and the elliptic operator $\Delta_{u}$. Their explicit analytical forms are provided in Eqs. (2)-(7), where the structural properties of convexity, coercivity, and ellipticity are established in detail. These constructions serve as the geometric and variational foundation of the Finsler-Ginzburg-Landau model under control translation, establishing a consistent link between anisotropic geometry and control theory.
@@
Building on this foundation, the paper develops the analytical framework of the control-translated Tonelli-Finsler energy and its asymptotic behavior. Through a detailed $\Gamma$-convergence analysis as $\varepsilon\to 0$, the renormalized energy $W_{u}$ is derived, governing vortex interactions in the presence of control and describing how $u(x)$ modifies both the Green kernel and the self-interaction potential. The final sections establish a variational and dynamical formulation, including the Euler-Lagrange system, the Hamiltonian structure of vortex motion, and the associated optimal control problem minimizing $E_{u}$.
@@
The results presented here extend the geometric program initiated in [1, 2] by embedding control theory directly into the convex-geometric formulation of the Finsler-Ginzburg-Landau model. The novelty of this work lies in the introduction of a control-translated Tonelli structure that couples Finsler geometry, convex analysis, and optimal control in a unified variational framework. This synthesis between Tonelli dynamics, Finsler geometry, and control analysis yields a mathematically consistent and geometrically interpretable mechanism for the stabilization, manipulation, and optimization of vortex structures in anisotropic superconducting media.
2 Control-Translated Tonelli Structure and Deformed Energy Functional
@@
remains smooth, strictly convex, and superlinear in $y$. Hence, it belongs to the class of Tonelli Lagrangians, that is, $C^{2}$ functions $L:TM\to\mathbb{R}$ which are strictly convex and superlinear in the fiber variable (cf. [7]). Tonelli Lagrangians generalize the analytic structure of Finsler geometry without assuming homogeneity, while preserving Legendre duality and Hamiltonian regularity. They provide a natural framework for control-dependent deformations where translation in the fiber variable encodes a geometric control field.
Let $(M,F)$ be a smooth compact manifold endowed with a strongly convex Finsler function $F:TM\to[0,\infty)$. Denote by $F^{*}(x,\xi)$ its convex dual. For a smooth control field $u\in C^{1}(M,TM)$ we define the control-translated Tonelli Lagrangian $\phi_{u}$ as in (2), and its convex conjugate on $T_{x}^{*}M$ by
@@
Let $\psi:M\to\mathbb{C}$ denote the order parameter and $A\in\Omega^{1}(M)$ a magnetic potential. Define the covariant derivative $D_{A}\psi=(d-iA)\psi$. The energy functional induced by the control-translated Tonelli structure is
@@
where $d\mu$ denotes the Busemann-Hausdorff volume form associated with the base Finsler structure $F$. Since the translation $y\mapsto y-u(x)$ preserves the volume element on each tangent fiber, $d\mu$ remains invariant under the control deformation. This choice ensures that the energy is geometrically well-defined and consistent with the underlying Tonelli structure (cf. [3]).
@@
Substituting (4) with $\xi=D_{A}\psi$ into (2) yields (9). The first term controls $\|D_{A}\psi\|_{L^{2}}$ by equivalence of Finsler and background norms. Using the Cauchy-Schwarz and Young inequalities,
@@
The Euler-Lagrange equations corresponding to $\mathcal{E}_{u}$ capture how the control interacts with the gauge and order fields through the geometric translation.
@@
Let $u_{k}\to u$ in $L^{\infty}(M;TM)$. Then $\mathcal{E}_{u_{k}}$ $\Gamma$-converges to $\mathcal{E}_{u}$ with respect to weak convergence in $H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$. Consequently, minimizers and gradient flows depend continuously on $u$ (cf. [6]).
@@
From (9), the $u$-dependence is affine and continuous in $L^{\infty}$. The $\Gamma$-liminf inequality follows from convexity and weak lower semicontinuity of $F^{*}(x,\cdot)^{2}$. A recovery sequence is given by $(\psi_{k},A_{k})=(\psi,A)$ because $\int_{M}\langle D_{A}\psi,u_{k}-u\rangle\,d\mu\to 0$. Therefore $\mathcal{E}_{u_{k}}\to\mathcal{E}_{u}$ in the $\Gamma$-sense. ∎
The functional $\mathcal{E}_{u}$ thus arises rigorously from the control-translated Tonelli Lagrangian $\phi_{u}$. The additional term $\langle D_{A}\psi,u\rangle$ is not a phenomenological assumption but an exact analytical consequence of convex duality. In this setting, control acts geometrically as a translation in the fiber variable, providing a mathematically consistent mechanism coupling geometry, control, and anisotropic superconductivity.
3 Stability and Renormalized Energy in Control-Translated Tonelli Structures
In this section we examine the asymptotic regime of the control-translated Tonelli Ginzburg-Landau functional $\mathcal{E}_{u}$ introduced previously, focusing on the emergence of renormalized energy, vortex concentration, and stability of critical configurations under control perturbations. The analysis combines the convex-analytic structure of $\phi_{u}^{*}$ with classical $\Gamma$-convergence arguments and vortex theory in anisotropic settings [4, 13].
Let $M$ be a compact oriented 2-dimensional manifold equipped with the Tonelli structure $(F_{u},F_{u}^{*})$ as defined in Section 2. For $\varepsilon>0$ sufficiently small, consider configurations $(\psi,A)\in H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$ minimizing
@@ 3 Stability and Renormalized Energy in Control-Translated Tonelli Structures
The measure $d\mu$ is the Busemann-Hausdorff volume of the base Finsler structure $F$, invariant under the translation $y\mapsto y-u(x)$ as explained before. Denote by $j_{A}=\Im(\bar{\psi}D_{A}\psi)$ the supercurrent and by $\rho=|\psi|^{2}$ the density.
@@
By the compactness and energy quantization principles of Bethuel-Brezis-Hélein [4] and the anisotropic ball construction and Jacobian estimates of Sandier-Serfaty [13], one deduces $\rho_{\varepsilon}\to 1$ in $L^{1}(M)$ and hence in measure. Writing $\psi_{\varepsilon}=\rho_{\varepsilon}^{1/2}e^{i\theta_{\varepsilon}}$, the current decomposes as $j_{A_{\varepsilon}}=\rho_{\varepsilon}(D\theta_{\varepsilon}-A_{\varepsilon})$. The coarea formula and the construction of vortex balls as in [4, 13] yield a finite family of disjoint balls $\{B(a_{i}^{\varepsilon},r_{i}^{\varepsilon})\}$ covering the defect set with $\sum_{i}r_{i}^{\varepsilon}\to 0$ and
@@
normalized by $\int_{M}G_{u}(x,y)\,d\mu(x)=0$. By the ellipticity of $\Delta_{u}$ (Theorem 2.2) and compactness of $M$, $G_{u}$ exists, is symmetric, and belongs to $C^{2,\alpha}_{\mathrm{loc}}(M\times M\setminus\mathrm{diag})$; this follows from Lax-Milgram, Fredholm alternative, and Schaüder theory [5], with the nonlinearity absorbed in the definition of $\Delta_{u}$ via the convex dual $\phi_{u}^{*}$ and the uniform positive definiteness [11].
@@
The function $\Phi_{u}$ encodes the effect of control translation on the self-interaction potential.
@@
Here we used the defining relation $\Delta_{u}G_{u}(\cdot,y)=\delta_{y}-1/\mu(M)$ and the fact that $\mathrm{curl}\,\mathcal{J}_{\varepsilon}=2\pi\sum_{i}d_{i}\delta_{a_{i}^{\varepsilon}}$ outside core balls. Passing to the limit along the convergent subsequence for $\nu_{\varepsilon}$ gives (16)-(17). The specific affine contribution (18) comes from the term $\int\langle\mathcal{J}_{\varepsilon},u\rangle d\mu$ via the identity
@@
which follows from the Green representation and Schauder theory for uniformly elliptic operators (see [5], and cf. [4] for the GL context]).
@@
In the limit $\varepsilon\to 0$, the supercurrent $j_{A}$ converges weakly to a measure-valued current $J_{u}$ satisfying $\mathrm{curl}\,J_{u}=2\pi\sum_{i}d_{i}\delta_{a_{i}}$ and
@@
The positions $a_{i}$ are local minimizers of $\mathcal{W}_{u}$, and under small perturbations of $u$ they vary smoothly according to
@@
Suppose $u$ is a smooth control field with small $\|u\|_{C^{1}}$. Then the equilibrium configuration of vortices minimizing $\mathcal{W}_{u}$ differs from the unperturbed lattice of the isotropic Finsler-Ginzburg-Landau model by an affine deformation of order $O(\|u\|_{C^{1}})$. Moreover, the corresponding minimal energy satisfies
@@
4 Variational Control and Vortex Dynamics in Tonelli-Finsler Geometry
The variational formulation developed in the previous sections naturally induces a dynamical framework where the control field $u(x)$ interacts with the geometric flow of vortices. We now derive and analyse the Euler-Lagrange equations of the control-translated Tonelli-Finsler energy $\mathcal{E}_{u}$ and describe the induced dynamics of vortex centers in the anisotropic setting.